OpenStudy (anonymous):

Is 2 the slope of (2, -3) and (6, 5)?

OpenStudy (anonymous):

I got y = 2x + -7 for the slope intercept form, can someone verify this for me?

jimthompson5910 (jim_thompson5910):

y = 2x - 7 is the equation that goes through those two points, nice work

jimthompson5910 (jim_thompson5910):

Check: y = 2x - 7 -3 = 2(2) - 7 ... plug in x = 2 and y = -3 -3 = 4 - 7 -3 = -3 So (2,-3) is on the line y = 2x - 7

jimthompson5910 (jim_thompson5910):

Check: y = 2x - 7 5 = 2(6) - 7 5 = 12 - 7 5 = 5 and (6,5) is also on the line y = 2x - 7

jimthompson5910 (jim_thompson5910):

Both points are on the line y = 2x-7, so that fully confirms the answer

OpenStudy (anonymous):

Can you help me with a few more questions?

OpenStudy (anonymous):

Write the equation of the line given the slope and the y-intercept. Type your answer in the box without skipping any spaces. m = 2, y-intercept (0, -9) I got y=2x+-9

jimthompson5910 (jim_thompson5910):

which is y = 2x - 9, nice job

OpenStudy (anonymous):

Finish writing the equation of the line y=1/4 x+?____, given the points (-12, 0) and (4, 4)

OpenStudy (anonymous):

I have no idea how to do that :(

jimthompson5910 (jim_thompson5910):

pick any point you want of the two, which one do you pick?

OpenStudy (anonymous):

-12

jimthompson5910 (jim_thompson5910):

the first one you meant?

jimthompson5910 (jim_thompson5910):

(-12, 0) right?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

ok (-12,0) means x = -12 and y = 0

jimthompson5910 (jim_thompson5910):

y=1/4 x+ b 0 = (1/4)(-12)+ b 0 = (1/4)(-12/1) + b 0 = -12/4 + b 0 = -3 + b .. ... b = ???

OpenStudy (anonymous):

3?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

y=1/4 x+?____ turns into y=1/4 x+3

OpenStudy (anonymous):

My last question...

OpenStudy (anonymous):

If a line has a slope of -2/5 then would a perpendicular line have a slope of 2/5?

jimthompson5910 (jim_thompson5910):

flip the fraction and flip the sign

jimthompson5910 (jim_thompson5910):

-2/5 ---> -5/2 ----> 5/2

jimthompson5910 (jim_thompson5910):

original slope: -2/5 perpendicular slope: 5/2

OpenStudy (anonymous):

Ohhh, ok then. I get it now

jimthompson5910 (jim_thompson5910):

the goal is the two slopes multiply to -1

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Well thank you so much for all you're help! :D

jimthompson5910 (jim_thompson5910):

you're welcome